inassamira2005:

Which equation represents y = x2 + 12x + 27 in vertex form? y = (x + 6)2 − 9 y = (x + 6)2 + 21 y = (x + 6)2 + 9 y = (x + 6)2 + 27

mhanifa:

Do you know how to complete a square?

inassamira2005:

I've heard of it, never used it in practice

mhanifa:

There are two ways: 1. Complete a square 2. Open parenthesis, simplify and compare

mhanifa:

@inassamira2005 wrote:
I've heard of it, never used it in practice
The it would be easier to go the second way. Use this identity for each option: (a + b)^2 = a^2 + 2ab + b^2

mhanifa:

Can you do that?

inassamira2005:

i can try, let me do that

mhanifa:

@inassamira2005 wrote:
i can try, let me do that
Sure

mhanifa:

How are you, need help?

inassamira2005:

ok would it be the last option because the product of the roots (-9,-3?) is 27?

mhanifa:

Apply (a + b)^2 = a^2 + 2ab + b^2 to (x + 6)^2 What you get?

mhanifa:

Replace a with x and b with 6. What is (x + 6)^2 = ...

inassamira2005:

(x + 6)^2= x^2 + 2x6 + 6^2 how would i do the 2x6 part?

mhanifa:

2 times 6 is ... and 6 times 6 is ...

inassamira2005:

12 and 36.. i still don't understand where this is all going..?

mhanifa:

(x + 6)^2= x^2 + 12x + 36 Now subtract 9 for the first option, what you get?

inassamira2005:

how did we get to subtracting 9? and subtract it from which number(s)

mhanifa:

y = (x + 6)^2 − 9 opened the parenthesis and got: y = x^2 + 12x + 36 - 9 Can you simplify?

inassamira2005:

x^2+12x+27

mhanifa:

Now you need to compare this with the given, are they same or different?

inassamira2005:

different..

mhanifa:

They are same and this option is your answer ... (x^2 and x2 is same in terms of your question)

inassamira2005:

ok ok so it would be the first option right? that is the on that is showing the same thing that we just did

mhanifa:

@inassamira2005 wrote:
ok ok so it would be the first option right? that is the on that is showing the same thing that we just did
Yes

inassamira2005:

thank you so much!

mhanifa:

No problem